I need to figure out this problem using substitution it's a word problem I would appreciate if let statements were used:
The local preschool ordered all new bicycles and tricycle for the new school year. Each bike,trike is shipped in a separate box. If there are 16 boxes of bicycles and tricycle total, and 45 wheels total, how many trikes were ordered?

Answer by macston(5194)

Let B=bicycles;Let T=tricycles
B+T=16
B=16-T
2B+3T=45 Substitute for B.Note: This equation is number of wheels time units.
2(16-T)+3T=45
32-2T+3T=45 Subtract 32 from each side.
T=13 ANSWER: 13 tricycles were ordered.
CHECK:
B=16-T=16-13=3 There were 3 bicycles
2B+3T=45
2(3)+3(13)=45
6+39=45
45=45